Hello, can you help me with this questions?

1). Assume that f(x) is one-to-one. Find f −1. f(x) = x3 − 11

2). Solve the equation for x. (Enter your answers as a comma-separated list. Round your answers to four decimal places.)

7x = 9x − 1
Answer #1

1) \large f(x)=x^3-11

to find inverse put f(x) as y and find equation for x . Then change x to f-1(x) and y to x to get this

on applying f(x) as y ,it becomes

\large y=x^3-11
find equation for x

\large x^3=y+11

\large x=\sqrt[3]{y+11}

now on changing it becomes

\large \boldsymbol{f^{-1}(x)=\sqrt[3]{x+11}}

so this is answer

2)

\large 7^x= 9^{x-1}

take natural log on both side

\large ln(7^x)= ln(9^{x-1})

we know ln (ab) = b*ln(a) .

so

\large x*ln(7)= (x-1)ln(9)

mon rearranging this

\large \frac{(x-1)}{x}=\frac{ln(7)}{ln(9)}

\large 1-\frac{1}{x}=\frac{ln(7)}{ln(9)}

\large \frac{1}{x}=1-\frac{ln(7)}{ln(9)}

\large x=\frac{1}{1-\frac{ln(7)}{ln(9)}}=8.7429

so answer is 8.7429